Weekly LIPAc highlights (11/October-15/October)


The maintenance of injector in the vault was completed. Accelerator column alignment was checked and found to be in line with last year checks. Boron Nitride disks were found to be strongly damaged after half year operation: they have been replaced by new ones (after baking). The cause of damage (surface eroded and colored in black) needs to be identified. Spares Boron Nitride disks are available on-site.

The PPS test was performed on Tuesday and HV up to 100kV were applied with no issues. D+ plasma conditioning was resumed on Wednesday this week.

Radio Frequency (RF) System and RFQ cavity

The last activities before resume RF-RFQ modules were completed, namely the installation of the re-calibrated VSWR cards into the PSYS units and the installation of the Anode Boxes with new PCB circuits into RFQ module 1 and 4B. The start-up process of RFQ-RF modules started on Wednesday.

An accidental spill of water occurred from a 50 kW dummy load during a series of planned tests on the SSPA-A. To check the SSPA functionalities after the accident, tests procedures were prepared and tests carried out: the SSPA-A functionality is unaltered, with a nominal output power extracted from the SSPA modules. A damage was found in an electronic board of the peripheral control unit and spare procurement started.

On the RFQ manifolds in the vault, the check of water flow rates on the hundreds of individual channels equipped with flow regulators was completed and results confirmed that no immediate actions is needed.

Beam Transport

The checks and calibration of the HEBT magnet PS is almost completed.

The mechanical check of the MEBT bunchers has been completed.

Functional checks of mass flow controllers for HEBT vacuum pressure control is on-going with CIEMAT remote support.


A new radiation resistant RGA and a new rupture disk were installed in vacuum sector IV of the HEBT.

All cryopumps for RFQ and HEBT have been successfully restarted.


Newly calibrated HEBT FPMs (Fluorescence Profile Monitors) were installed. Functional check will be done though the remote session with CIEMAT.

Checks are ongoing on BD accelerometer cables with CIEMAT remote support.

Control system.

The integration of common chiller LCS into the CCS is on-going. The communication line between BD skid and HEBT LCS has been renewed.

In the current Integrated LIPAc schedule, it is planned to resume beam operation early November.

HEBT sector IV, new RGA radiation resistant + new rupture disk

Disclaimer: the information reported is not meant to be technically complete and doesn’t cover all the activities currently carried out on LIPAc.